# Quantum automata and languages of finite index thanks:

### v1

@@ Abstract
This paper is a continuation of a previous study conducted in [7, 8] on the so-called measure once finite quantum automata model introduced by Moore and Crutchfield in 2000. We investigate conditions assuring that, given a language recognized by such a device and a language generated by a context-free grammar of finite index or by a matrix context-free grammar, it is recursively decidable whether or not they have a nonempty intersection.
@@ 1 Introduction
In this paper, we investigate an issue of decidability for quantum (finite) automata. Quantum automata were introduced at the beginning of 2000’s in [23] as a new model of language recognizer (cf. [5]). Numerous publications have ever since compared their decision properties to those of other models and, in particular, to probabilistic finite automata. In this theoretical setting, we investigate the decidability of a problem for quantum automata introduced in [6, 7], named Intersection Problem. The starting point of that investigation is a result of Blondel, Jeandel, Koiran, and Portier which shows in [8] that, given a (“measure-once” in the sense of [23]) quantum automaton $\cal Q$, it is decidable whether the language recognized by $\cal Q$ has a non trivial intersection with the free monoid generated by the input alphabet of $\cal Q$ (Emptiness problem).
It is worth remarking that these problems are strictly related to some computational issues concerning matrices. Precisely, a crucial step in the construction of the decision procedure for both the Emptiness and Intersection Problem amounts to effectively compute the Zariski closure of a finitely generated group of matrices over $\mathbb{Q}$. For that task, an appropriate algorithm has been developed in [12] by Derksen, Jeandel, and Koiran.
Before coming back to the Intersection Problem, we find then convenient to recall some developments of the research on that aspect. Two striking results have been obtained in the recent past years in connection with Derksen et al. algorithm. In 2022 Nosan, Pouly, Schmitz, Shirmohammadi, and Worrell, following an alternative approach for such computation, have obtained in [25] a bound on the degree of the polynomials that define the Zariski closure of the set and showed that the latter can be computed in elementary time.
The second was obtained in 2023 by Hrushovski, Ouaknine, Pouly, and Worrel for the study of the effective computation of polynomial invariants for affine programs [17]. Precisely, it is exhibited an algorithm that, given a finite set of square matrices of the same size with coefficients in $\mathbb{Q}$, computes the Zariski closure of the semigroup generated by the set.
It should be pointed out that the decidability issues for matrix semigroups is a challenge problem. Even in small size, a well-known result by Paterson shows that the problem of freeness and that of membership are both undecidable for matrix semigroups over $\mathbb{Z}$ of size $3$ (cf [9]). In contrast, in 2017, a surprising result of Potapov and Semukhin [28] shows that the membership is decidable for non-singular matrix integer semigroups of size $2$. We now come back to the Intersection Problem and give a formal presentation of our work. A finite quantum automaton (in the sense of [23]) is a quadruple
@@
where |||||$\cdot$||| | ⋅ | | denotes the Euclidean norm of vectors. Let ${\cal L}$ be a family of effectively defined languages. The problem we tackle is thus the following:
@@
As already pointed out, it has been proven in [8] the decidability of the problem in the specific case $L$ $=\Sigma^{*}$. This result contrasts with the corresponding problems (strict and non strict, respectively) for probabilistic finite automata. Indeed for this class of automata, the Emptiness problems w.r.t. $|{\cal Q}_{>}|$ and $|{\cal Q}_{\geq}|$ are both undecidable (see [27], Thm 6.17). Also it contrasts with the undecidability of the Emptiness problem of $|{\cal Q}_{\geq}|$ for quantum automata [8]. It is also proven in [21] that both the problems (strict and non strict, respectively) remain undecidable for the “measure many” model of quantum automata, a model not computationally equivalent to the “measure once”, introduced by Kondacs and Watrous in [22].
@@
– where $M$ is an arbitrary matrix in $\mathbb{R}^{n\times n}$, and $s,P$ are components of (1) – is continuous, it is sufficient to prove that, for all matrices $M$ in the Euclidean closure ${\mathbf{C}l}(\varphi(L))$ of $\varphi(L)$, the condition $||sMP||\leq\lambda$ holds (see Section 3). Such condition is then effectively tested by running, in parallel, two semialgorithms. The first tests the inclusion $\varphi(L)\subseteq|{\cal Q}_{>}|$ and it is readily seen to be semidecidable. The second, which is the main bulk, is to prove that (2) is semidecidable.
The semidecidability of the latter procedure is achieved by expressing the property (2) in first-order logic of the field of reals, which, in turn, consists to effectively compute a representation of ${\mathbf{C}l}(\varphi(L))$ in term of semialgebraic set (a more general property than algebraic, which allows to cope with the fact that algebraic sets are not closed w.r.t. the operation of product of sets). Afterwards, one applies on the constructed formula the Tarski-Seidenberg quantifier elimination to verify whether (2) holds true or not.
It turns out that the semialgebraicity of ${\mathbf{C}l}(\varphi(L))$, and its effective computation (as a semialgebraic set), are reduced to those of an algebro-combinatorial object that is associated with the grammar $G$ generating $L$, and with every variable $A$ of $G$: the monoid of cycles of $A$. Such monoid, denoted $M_{A}$, corresponds (up to a technical detail) to the matrix image, under taking the morphism $\varphi$ of (1), of the language of cycles in $G$ associated with $A$ (see Eq. (21)).
The reduction process mentioned above is made possible by resorting to two ingredients: the fact that the Euclidean closure ${\mathbf{C}l}(M_{A})$ is an algebraic set (because of a deep result of Algebraic Geometry, namely Theorem 3) and because of a suitably defined effective structuring of the computations, i.e., the derivations, of $G$ (Sections 4.1, 4.2 and 6.1). In this theoretical setting, [7] and this work provides the following picture of the Intersection Problem:
@@
The new results of the paper are the following. In Section 6.1, we extend the main contribution of [7] – i.e., the statement of decidability of the Intersection Problem for linear context-free languages and for bounded semilinear languages – to a class of matrix context-free grammar of finite index. Precisely, whenever the quantum automaton is rational, such result (Proposition 10) allows, on one hand, to recover both the previous two cases as immediate corollaries of such result. On the other, it shows the decidability of the problem for languages of exponential growth that are not context-free. The proof is based upon a combinatorial decomposition of these matrix grammars proposed by Choffrut, Varricchio and the second author in [10] for the study of rational relations. In particular, this decomposition permits a refinement of the algebro-combinatorial machinery developed in [7] for the manipulation of the monoids of cycles of these grammars.
@@
We prove that, whatever is the length $k$ of the composition (3) for $G$, if the Zariski closure of the monoids of cycles of the linear grammars of the lowest level ${\cal G}_{k}$ of (3), are (algebraic) irreducible, then one can effectively computes the Euclidean closure ${\mathbf{C}l}(\varphi(L))$ of $\varphi(L)$ (Proposition 11, Corollary 5).
Extensions of such result to broader families of finite index context-free languages seem related to possible extensions of the algorithms [12, 17, 25] (cf Section 7, [$\alpha$]).
The general case of context-free languages is still open and it would be even possible that the problem is undecidable (cf Section 7, [$\beta$]).
@@ 2.2 Algebraic and semialgebraic sets
The set ${\cal A}$ is said to be effective algebraic if the polynomials of $\cal P$ can be computed by an effective procedure (algorithm).
Note that by the Hilbert basis Theorem, one may assume that the set $\cal P$ is finite. Even more, since we are dealing with algebraic sets over $\mathbb{R}$, $\cal P$ can be reduced to a singleton since (5) can be assumed to be to the equation
@@
The set ${\cal A}$ is said to be effective semialgebraic if the formula defining ${\cal A}$ can be effectively computed.
We now specify such definitions to square matrices.
@@
We now combine the notions of zero sets and of topology. In the following we rephrase two results of [8] by emphasizing the main features that serve our purpose. Given a subset $E$ of a group, we denote by $\langle E\rangle$ and by $E^{*}$ the subgroup and the submonoid it generates, respectively.
@@
The main consequence of the next theorem is that the topological closure of a monoid of orthogonal matrices is algebraic ([26], Ch. 3, Sec. 4, Theorem $5$).
@@ 2.4 Algebraic Irreducible sets
Let us consider the topological space obtained by equipping $\mathbb{R}^{m}$ with the Zariski topology. Given a non-empty subset $\cal Z$ of $\mathbb{R}^{m}$, $\cal Z$ is said to be irreducible if $\cal Z$ cannot be written as the union of two closed, i.e. algebraic, subsets of $\cal Z$, distinct from $\emptyset$ and $\cal Z$ itself. It can be proven that a non-empty algebraic subset $\cal Z$ of $\mathbb{R}^{m}$ is irreducible if and only if its vanishing ideal is a prime ideal (cf [13], Prop. 1.1.12). Irreducible sets will play a role in this theoretical setting (cf Sec. 6.2). To this purpose, we recall some basic results. The next notation is useful.
@@
Given non-empty algebraic sets ${\cal V}\subseteq\mathbb{R}^{n}$ and ${\cal W}\subseteq\mathbb{R}^{m}$, a map $f:{\cal V}\rightarrow~{}{\cal W}$ is said to be regular444 By following a well-established vocabulary in Automata Theory, the term regular will be also used in Sec. 4 to name the family of subsets of monoids defined by finite automata (or a morphism of algebraic sets), if there exist $p_{1},\dots,p_{m}\in\mathbb{R}[x_{1},.\ldots,x_{n}]$ (where the $x_{i}$ are indeterminates) such that, for all ${\bf x}=(x_{1},\ldots,n_{n})$ in $\cal V$
@@ 4 Preliminaries on context-free languages
A context-free grammar $G$ is a quadruple $\langle V,\Sigma,P,S\rangle$ where $\Sigma$ is the alphabet of terminal symbols, $V$ is the set of variables or nonterminal symbols, $P$ is the set of rules or productions, and $S$ is the axiom of the grammar. A word over the alphabet $\Sigma$ is called terminal. As usual, the nonterminal symbols are denoted by uppercase letters as, for instance, $A$, $B$. An arbitrary rule of the grammar is written as $A\rightarrow\alpha$. The derivation relation of $G$ is denoted by $\mathop{\Rightarrow}^{*}$. The language generated by $G$ is the subset of terminal words $\{w\in\Sigma^{*}:S\displaystyle\mathop{\Rightarrow}^{*}w\}$ denoted by $L(G)$.
@@ 5 Context-free languages and semialgebraic sets
$M_{A}=\{\varphi(u_{1})\oplus\varphi(u_{2})^{\mbox{T}}\mid(u_{1},u_{2})\in C_{A}\},$ (21)
where $\varphi(u_{2})^{\mbox{T}}$ denotes the transpose of $\varphi(u_{2})$. $M_{A}$ will be called the monoid of cycles of $A$ since the following holds.
@@
Observe first that $M_{A}$ is a subsemigroup of the group $O_{n}\oplus O_{n}.$ Indeed, if $\varphi(u_{1})\oplus\varphi(u_{2})^{\mbox{T}}$ and $\varphi(v_{1})\oplus\varphi(v_{2})^{\mbox{T}}$ are in $M_{A}$ then we have
$(\varphi(u_{1})\oplus\varphi(u_{2})^{\mbox{T}})(\varphi(v_{1})\oplus\varphi(v_{2})^{\mbox{T}})=\varphi(u_{1}v_{1})\oplus\varphi(u_{2})^{\mbox{T}}\varphi(v_{2})^{\mbox{T}},$ (22)
@@
$\varphi(u_{1}v_{1})\oplus\varphi(v_{2}u_{2})^{\mbox{T}}$ (23)
equals (22) for $\varphi(v_{2}u_{2})^{\mbox{T}}=\varphi(u_{2})^{\mbox{T}}\varphi(v_{2})^{\mbox{T}}.$ Since $(\varepsilon,\varepsilon)\in M_{A}$, then $M_{A}$ is a monoid. The second statement comes from Theorems 2 and 3. ∎∎
@@
are algebraic. If $\mathop{\mathrm{Card}}\nolimits(V)=1$, i.e., the set of nonterminal symbols is reduced to $S$ then $L$ is reduced to $C_{S}\diamond L^{\prime}(G)$ and $L^{\prime}(G)$ is finite. We may further assume that there is a sole terminal rule $S\rightarrow w$, so that $L^{\prime}(G)=\{w\}$. By Theorem 1 one has [^6] By Proposition 1, if ${\mathbf{C}l}(M_{S})$ is algebraic, then $\{X\oplus Y^{\mbox{T}}:X\oplus Y\in{\mathbf{C}l}(M_{S})\}$ is so
${\mathbf{C}l}(\varphi(L))=\{X\varphi(w)Y^{\mbox{T}}\mid X\oplus Y\oplus\{\varphi(w)\}\in{\mathbf{C}l}(M_{S}\oplus\varphi(w))\}$ (26)
@@
${\mathbf{C}l}(\varphi(C_{S})\diamond\varphi(L^{\prime\prime}))=\{XZY^{\mbox{T}}\mid X\oplus Y\oplus Z\in{\mathbf{C}l}(M_{S})\oplus{\mathbf{C}l}(\varphi(L^{\prime\prime}))\}=$
$\{XZY^{\mbox{T}}\mid X\oplus Y\oplus Z\in{\mathbf{C}l}(M_{S})\oplus{\mathbf{C}l}(\varphi(w_{1})\varphi(L_{A_{1}})\cdots\varphi(w_{\ell})\varphi(L_{A_{\ell}})\varphi(w_{\ell+1}))\}$

### v2

@@ Abstract
This paper continues the study of measure-once finite quantum automata building on work by Bertoni et al. and Blondel et al. We investigate conditions ensuring that, given a language recognized by such a device and a language generated by a context-free grammar of finite index or by a matrix context-free grammar, it is decidable whether or not they have a nonempty intersection.
@@ 1 Introduction
In this paper, we investigate an issue of decidability for quantum (finite) automata. Quantum automata were introduced at the beginning of 2000’s in [23] as a new model of language recognizer (cf. [5]). Numerous publications have ever since compared their decision properties to those of other models and, in particular, to probabilistic finite automata. In this theoretical setting, we investigate the decidability of a problem for quantum automata introduced in [6, 7], named Intersection Problem. The starting point of that investigation is a result of Blondel, Jeandel, Koiran, and Portier who showed in [8] that, given a (“measure-once” in the sense of [23]) quantum automaton $\cal Q$, it is decidable whether the language recognized by $\cal Q$ has a non trivial intersection with the free monoid generated by the input alphabet of $\cal Q$ (Emptiness problem). It is worth remarking that these problems are strictly related to some computational issues concerning matrices. Precisely, a crucial step in the construction of the decision procedure for both the Emptiness and Intersection Problem amounts to effectively compute the Zariski closure of a finitely generated group of matrices over $\mathbb{Q}$. For that task, an appropriate algorithm has been developed in [12] by Derksen, Jeandel, and Koiran.
Before coming back to the Intersection Problem, we find then convenient to recall some developments of the research on that aspect. Two striking results have been obtained recently in connection with the algorithm mentioned above. In 2022 Nosan, Pouly, Schmitz, Shirmohammadi, and Worrell, following an alternative approach for such computation, have obtained in [25] a bound on the degree of the polynomials that define the Zariski closure of the set and showed that the latter can be computed in elementary time.
The second was obtained in 2023 by Hrushovski, Ouaknine, Pouly, and Worrell for the study of the effective computation of polynomial invariants for affine programs [17]. Precisely, it is exhibited an algorithm that, given a finite set of square matrices of the same size with coefficients in $\mathbb{Q}$, computes the Zariski closure of the semigroup generated by the set.
It should be pointed out that the decidability issues for matrix semigroups is a challenging problem. Even in small size, a well-known result by Paterson shows that the problem of freeness and that of membership are both undecidable for matrix semigroups over $\mathbb{Z}$ of size $3$ (see [9]). In contrast, in 2017, a surprising result of Potapov and Semukhin [28] shows that the membership is decidable for non-singular matrix integer semigroups of size $2$.
We now come back to the Intersection Problem and give a formal presentation of our work. A finite quantum automaton (in the sense of [23]) is a quadruple
@@
where |||||$\cdot$||| | ⋅ | | denotes the Euclidean norm of vectors. Let ${\cal L}$ be a family of effectively defined languages. The problem we tackle is the following:
@@
As already pointed out, the decidability of the above-mentioned problem has been proven in [8] in the specific case $L$ $=\Sigma^{*}$. This result contrasts with the corresponding problems (strict and non strict, respectively) for probabilistic finite automata. Indeed for this class of automata, the Emptiness problems w.r.t. $|{\cal Q}_{>}|$ and $|{\cal Q}_{\geq}|$ are both undecidable (see [27], Thm 6.17). Also it contrasts with the undecidability of the Emptiness problem of $|{\cal Q}_{\geq}|$ for quantum automata [8]. It is also proven in [21] that both the problems (strict and non strict, respectively) remain undecidable for the “measure many” model of quantum automata, a model not computationally equivalent to the “measure once”, introduced by Kondacs and Watrous in [22].
@@
– where $M$ is an arbitrary matrix in $\mathbb{R}^{n\times n}$, and $s,P$ are components of (1) – is continuous, it is sufficient to prove that, for all matrices $M$ in the Euclidean closure ${\mathbf{C}l}(\varphi(L))$ of $\varphi(L)$, the condition $||sMP||\leq\lambda$ holds (see Section 3). Such condition is then effectively tested by running, in parallel, two semialgorithms. The first tests the inclusion $\varphi(L)\subseteq|{\cal Q}_{>}|$ and it is readily seen to be semidecidable. The second, which is the main crux, is to prove that (2) is semidecidable.
The semidecidability of the latter procedure is achieved by expressing the property (2) in first-order logic of the field of reals, which, in turn, consists to effectively compute a representation of ${\mathbf{C}l}(\varphi(L))$ in term of a semialgebraic set (a more general property than algebraic, which allows one to cope with the fact that algebraic sets are not closed w.r.t. the operation of product of sets). Afterwards, one applies on the constructed formula the Tarski-Seidenberg quantifier elimination to verify whether (2) holds true or not.
It turns out that, in the case the language $L$ is context-free, the semialgebraicity of ${\mathbf{C}l}(\varphi(L))$, and its effective computation (as a semialgebraic set), are reduced to those of an algebro-combinatorial object that is associated with the grammar $G$ generating $L$, and with every variable $A$ of $G$: the monoid of cycles of $A$. Such monoid, denoted $M_{A}$, corresponds (up to a technical detail) to the matrix image, under taking the morphism $\varphi$ of (1), of the language of cycles in $G$ associated with $A$ (see Eq. (21)).
The reduction process mentioned above is made possible by resorting to two ingredients: the fact that the Euclidean closure ${\mathbf{C}l}(M_{A})$ is an algebraic set (because of a deep result of Algebraic Geometry, namely Theorem 3) and because of a suitably defined effective structuring of the computations, i.e., the derivations, of $G$ (Sections 4.1, 4.2 and 6.1). In this theoretical setting, the work in [7] and this work provide the following picture of the Intersection Problem:
@@
The new results of the paper are the following. In Section 6.1, we extend the main contribution of [7] – i.e., the statement of decidability of the Intersection Problem for linear context-free languages and for bounded semilinear languages – to a class of matrix context-free grammar of finite index. Precisely, whenever the quantum automaton is rational, such result (Proposition 10) allows one, on one hand, to recover both the previous two cases as immediate corollaries of such result. On the other, it shows the decidability of the problem for languages of exponential growth that are not context-free. In the proof a suitable decomposition of the derivations in these grammars is used. Such a decomposition is based upon some combinatorial tools proposed by Choffrut, Varricchio and the second author in [10] for the study of rational relations. In particular, this decomposition permits a refinement of the machinery developed in [7] for the manipulation of the monoids of cycles of these grammars.
@@
We prove that, regardless of the length $k$ of the composition (3) for $G$, if the Zariski closure of the monoids of cycles of the linear grammars of the lowest level ${\cal G}_{k}$ of (3), are (algebraic) irreducible, then one can effectively compute the Euclidean closure ${\mathbf{C}l}(\varphi(L))$ of $\varphi(L)$ (Proposition 11, Corollary 5).
Extensions of such result to broader families of finite index context-free languages seem related to possible extensions of the algorithms [12, 17, 25] (cf Section 7, [$\alpha$]). The general case of context-free languages is still open and it would be even possible that the problem is undecidable (cf Section 7, [$\beta$]).
@@ 2.2 Algebraic and semialgebraic sets
Note that by the Hilbert’s basis theorem (see [13], Ch. 1, Sec. 1, Theorem $1$), one may assume that the set $\cal P$ is finite. Even more, since we are dealing with algebraic sets over $\mathbb{R}$, $\cal P$ can be reduced to a singleton since (5) can be assumed to be to the equation
@@
We now instantiate the definitions above to the setting of square matrices.
@@
Notational convention. In connection with the $(L,{\cal Q})$ Intersection Problem, we will adopt the following terminology. A set ${\cal A}$ of matrices will be called effective algebraic (resp., effective semialgebraic) if the polynomials (resp., the formula) defining ${\cal A}$ can be algorithmically computed from the input of the problem, i.e., an effectively defined language $L$ and a finite quantum automaton $Q$.
We now combine the notions of zero sets and of topology. In the following we rephrase two results of [8] by emphasizing the main features that serve our purpose. Given a subset $E$ of a group, we denote by $\langle E\rangle$ and by $E^{*}$ the subgroup and the submonoid generated by $E$, respectively.
@@
The main consequence of the next theorem is that the topological closure of a monoid of orthogonal matrices is algebraic ([26], Ch. 3, Sec. 4, Theorem $5$, [7], Theorem $7$).
@@ 2.4 Algebraic Irreducible sets
Let us consider the topological space obtained by equipping $\mathbb{R}^{m}$ with the Zariski topology (see [13], Ch. $1$). Given a non-empty subset $\cal Z$ of $\mathbb{R}^{m}$, $\cal Z$ is said to be irreducible if $\cal Z$ cannot be written as the union of two closed, i.e. algebraic, subsets of $\cal Z$, distinct from $\emptyset$ and $\cal Z$ itself. It can be proven that a non-empty algebraic subset $\cal Z$ of $\mathbb{R}^{m}$ is irreducible if and only if its vanishing ideal is a prime ideal (cf [13], Prop. 1.1.12). Such sets play a role in the theoretical developments presented later (see Sec. 6.2). Therefore, we recall a definition and some results. The next notation is useful.
@@
Given non-empty algebraic sets ${\cal V}\subseteq\mathbb{R}^{n}$ and ${\cal W}\subseteq\mathbb{R}^{m}$, a map $f:{\cal V}\rightarrow~{}{\cal W}$ is said to be regular111 By following a well-established vocabulary in Automata Theory, the term regular will be also used in Sec. 4 to name the family of subsets of monoids defined by finite automata (or a morphism of algebraic sets), if there exist $p_{1},\dots,p_{m}\in\mathbb{R}[x_{1},.\ldots,x_{n}]$ (where the $x_{i}$ are indeterminates) such that, for all ${\bf x}=(x_{1},\ldots,x_{n})$ in $\cal V$
@@ 4 Preliminaries on context-free languages
A context-free grammar $G$ is a quadruple $\langle V,\Sigma,P,S\rangle$ where $\Sigma$ is the alphabet of terminal symbols, $V$ is the set of variables or nonterminal symbols, $P$ is the set of rules or productions, and $S$ is the axiom of the grammar. A word over the alphabet $\Sigma$ is called terminal. As usual, the nonterminal symbols are denoted by uppercase letters as, for instance, $A$, $B$. An arbitrary rule of the grammar is written as $A\rightarrow\alpha$, where $\alpha$ is a finite sequence of variables and terminals. The derivation relation of $G$ is denoted by $\mathop{\Rightarrow}^{*}$. The language generated by $G$ is the subset of terminal words $\{w\in\Sigma^{*}:S\displaystyle\mathop{\Rightarrow}^{*}w\}$ denoted by $L(G)$.
@@ 5 Context-free languages and semialgebraic sets
$M_{A}=\{\varphi(u_{1})\oplus\varphi(u_{2})^{\scriptstyle T}\mid(u_{1},u_{2})\in C_{A}\},$ (21)
where $\varphi(u_{2})^{\scriptstyle T}$ denotes the transpose of $\varphi(u_{2})$. $M_{A}$ will be called the monoid of cycles of $A$ since the following holds.
@@
Observe first that $M_{A}$ is a subsemigroup of the group $O_{n}\oplus O_{n}.$ Indeed, if $\varphi(u_{1})\oplus\varphi(u_{2})^{\scriptstyle T}$ and $\varphi(v_{1})\oplus\varphi(v_{2})^{\scriptstyle T}$ are in $M_{A}$ then we have
$(\varphi(u_{1})\oplus\varphi(u_{2})^{\scriptstyle T})(\varphi(v_{1})\oplus\varphi(v_{2})^{\scriptstyle T})=\varphi(u_{1}v_{1})\oplus\varphi(u_{2})^{\scriptstyle T}\varphi(v_{2})^{\scriptstyle T},$ (22)
@@
$\varphi(u_{1}v_{1})\oplus\varphi(v_{2}u_{2})^{\scriptstyle T}$ (23)
equals (22) for $\varphi(v_{2}u_{2})^{\scriptstyle T}=\varphi(u_{2})^{\scriptstyle T}\varphi(v_{2})^{\scriptstyle T}.$ Since $(\varepsilon,\varepsilon)\in M_{A}$, then $M_{A}$ is a monoid. The second statement comes from Theorems 2 and 3. ∎∎
@@
are algebraic. If $\mathop{\mathrm{Card}}\nolimits(V)=1$, i.e., the set of nonterminal symbols is reduced to $S$ then $L$ is reduced to $C_{S}\diamond L^{\prime}(G)$ and $L^{\prime}(G)$ is finite. We may further assume that there is a sole terminal rule $S\rightarrow w$, so that $L^{\prime}(G)=\{w\}$. By Theorem 1 one has [^3] By Proposition 1, if ${\mathbf{C}l}(M_{S})$ is algebraic, then $\{X\oplus Y^{\scriptstyle T}:X\oplus Y\in{\mathbf{C}l}(M_{S})\}$ is so
${\mathbf{C}l}(\varphi(L))=\{X\varphi(w)Y^{\scriptstyle T}\mid X\oplus Y\oplus\{\varphi(w)\}\in{\mathbf{C}l}(M_{S}\oplus\varphi(w))\}$ (26)
@@
${\mathbf{C}l}(\varphi(C_{S})\diamond\varphi(L^{\prime\prime}))=\{XZY^{\scriptstyle T}\mid X\oplus Y\oplus Z\in{\mathbf{C}l}(M_{S})\oplus{\mathbf{C}l}(\varphi(L^{\prime\prime}))\}=$
$\{XZY^{\scriptstyle T}\mid X\oplus Y\oplus Z\in{\mathbf{C}l}(M_{S})\oplus{\mathbf{C}l}(\varphi(w_{1})\varphi(L_{A_{1}})\cdots\varphi(w_{\ell})\varphi(L_{A_{\ell}})\varphi(w_{\ell+1}))\}$